A passenger is standing inside a moving bus. The bus negotiates the u turn in the chandmari flyover at a speed v. If the radius of curvature of the u turn is R then find the angle through which the passenger leans away from the vertical
1
Expert's answer
2019-10-28T10:57:37-0400

We have:


"tan{\theta}=\frac{W}{F_c}"

"tan{\theta}=\frac{mgR}{mv^2}=\frac{gR}{v^2}"

The angle through which the passenger leans away from the vertical:


"\theta=\arctan{\frac{gR}{v^2}}"